# Faster shortest-path algorithms using the acyclic-connected tree

### v1

@@ 7 Example
Consider again Example 1.1 from the introduction. Running algorithm A-C_tree (Algorithm 1) gives us the A-C tree of $G$ in Figure 2(b). By Theorem 4.1, this A-C tree defines a minimal-width nesting decomposition of $G$ with width 3. This $G$ demonstrates that nesting decomposition can be quite non-trivial, even in strongly connected graphs. A small-width decomposition can be found because $G$ has an ‘almost acyclic’ structure.
The algorithm also identifies the nesting bottleneck in this structure, which is the strongly connected set $\{e,d\}$. Indeed, when computing shortest paths using Recursive Dijkstra’s (Algorithm 3), the maximum quueu length used in the algorithm is equal to $\max_{i}|K_{i}|=|\{e,d\}|=2$, where maximum is taken over all strongly connected components $K_{i}$ in the A-C tree. This is a significant improvement compared to naïvely running Dijkstra’s algorithm, with a single large queue of length $9$ containing all nodes.
@@ 8.1 Summary
In this paper, we introduced a graph decomposition called the A-C tree to provide parametrised complexity bounds for SSSP. The A-C tree breaks up a single-source graph $G$ into a recursively nested sequence of strongly connected components in topological order. We showed that the A-C tree is optimal in the sense that it achieves the nesting width $\operatorname{nw}(G)$, the minimal width of any nesting decomposition of $G$. We also presented a linear-time algorithm for constructing the A-C tree, which we used to design a SSSP algorithm called Recursive Dijkstra’a with time complexity $O(e+n\log(\operatorname{nw}(G)))$.
@@ 8.2 Future Work
Our results demonstrate that, even on this classical and fundamental problem, new graph ideas can still lead to new insight. Among other things, we highlight how despite the important progress on the complexity of the distance ordering problem (DO), the—arguably more natural—SSSP problem is open in many cases. The fact that a combination of simple linear-time graph algorithms can lead to non-trivial advances on this problem suggest that there is room for significant progress in future work. The direct follow-up questions are the following:
First, classifying the classes of graphs with bounded nesting width is important. In a sense, these graphs are ‘somewhat acyclic’ (generalising the DAG case), and as a result, SSSP can be solved in linear time. As the example in this paper illustrate, these classes of graphs are surprisingly rich, and are therefore worth classifying. Second, it would be interesting to understand the frequency that graphs with low nesting width occur in (sparse) random graphs. Thirdly, our approach is entirely structural and does not depend on the arc weights. This resembles universal optimality in the sense of [28]. Our results could be a first step towards building a universally optimal algorithm for SSSP, similar to the results of [28] for DO. Fourthly, and similarly, our independence from weights could be relevant in application domains where, for example, the weights can change over time but the graph topology does not. One motivating example is road networks, where traffic conditions change over time.

### v2

@@ 7 Example
Let us return to Figure 1 from the introduction. Running algorithm A-C_tree (Algorithm 1) gives us the A-C tree of $G$ shown in Figure 2(b). By Theorem 4.1, this A-C tree defines a minimal-width nesting decomposition of $G$ with width 3. The graph $G$ is strongly connected, but its A-C tree is quite non-trivial with a small nesting width. Intuitively, $G$ has an ‘almost acyclic’ structure.
The algorithm also identifies the nesting bottleneck in this structure, which is the strongly connected set $\{e,d\}$. Indeed, when computing shortest paths using Recursive Dijkstra (Algorithm 3), the maximum queue length used in the algorithm is equal to $\max_{i}|K_{i}|=|\{e,d\}|=2$, where maximum is taken over all strongly connected components $K_{i}$ in the A-C tree. This is a significant improvement compared to a naïve execution of Dijkstra’s algorithm, with a single large queue of length $9$ containing all nodes.
@@ 8.1 Summary
In this paper, we introduced a graph decomposition called the A-C tree to obtain beyond-worst-case time complexity for any SSSP algorithm. The A-C tree breaks up a single-source graph $G$ into a recursively nested sequence of strongly connected components in topological order. We showed that the A-C tree is optimal in the sense that it achieves the nesting width $\operatorname{nw}(G)$, the minimal width of any nesting decomposition of $G$. We also presented a linear-time algorithm for constructing the A-C tree. Using this linear-time algorithm, we can transform an SSSP algorithm by first computing the A-C tree, and then run the SSSP algorithm on the A-C tree, for reduced time complexity. Indeed, we showcased this transformation with state-of-the-art Dijkstra’s algorithm and the recent sparse graph algorithm from [19], with improved time complexities $O(m+n\log(\operatorname{nw}(G)))$ and $O(m\alpha(n)+m\log^{2/3}(\operatorname{nw}(G)))$, respectively. In particular, SSSP for graphs with bounded nesting width can be solved in linear time, an important step towards a universally optimal SSSP algorithm.
@@ 8.2 Future Work
Our results demonstrate that even on this classical and fundamental problem, new graph ideas can still lead to new insights. Among other things, we highlight how, despite the important progress on the complexity of DO, the arguably more natural SSSP problem is open in many cases. The fact that a combination of well-known linear-time graph algorithms can lead to non-trivial advances on this problem suggest that there is room for significant progress in future work. Direct follow-up questions are as follows.
First, classifying the classes of graphs with bounded nesting width is important. In a sense, these graphs are ‘somewhat acyclic’, generalising DAGs, and as a result, SSSP can be solved in linear time. As the example in this paper illustrate, these classes of graphs are surprisingly rich, and are therefore worth classifying. Second, it would be interesting to understand the frequency that graphs with low nesting width occur in (sparse) random graphs. Third, our approach is entirely structural and does not depend on the arc weights. This resembles universal optimality in the sense of [29]. Our results could be a first step towards building a universally optimal algorithm for SSSP, similar to the results of [29] for DO. Fourth, there are several related problems to SSSP, such as graphs with stochastic transitions between nodes, that could potentially apply our method for improved performance. Fifth, our independence from weights could also be relevant in application domains where, for example, the weights can change over time but the graph topology does not. One motivating example is road networks, where traffic conditions change over time.